OpenStudy (anonymous):

solve for m: F=GMm/r^2

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Fr^2=GMm ((Fr^2)/GM)=m

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

I actually couldn't get past the (r^2)F=GMm because I didn't know if (r^2)F= Fr^2

OpenStudy (anonymous):

yes they are the same or equal

OpenStudy (anonymous):

I get it. Thanks!

OpenStudy (anonymous):

either you multiply them from the left or to the right, they are the same, like 2x3=3x2

OpenStudy (anonymous):

welcome